Suppose a professor does 190000 J of work while metabolizing 490 kcal of food energy. What is the efficiency of the professor? How much energy in kcal will an athlete metabolize in doing the same work with an efficiency of 0.20?
1
Expert's answer
2018-03-07T11:41:07-0500
"efficiency=" "useful work" /"expended energy"
So, efficiency of the professor
η=(190000 J)/(490 kcal)=(190000 J)/(490×4184 J)=0.093=9.3%
The amount of energy that athlete will metabolize
190000/0.20=950000 J=227 kcal